OpenStudy (anonymous):

if f(x) = -x+4, which of the following is not true? The choices are: a. f(1) < f(0) b. f(1) > f(-1) c. f(-1) > f(0) d. f(0) < f(-1). My prof's answer was letter B. why is it so?

OpenStudy (anonymous):

f(x) = -x= 4? Did you type this correctly?

OpenStudy (anonymous):

Sorry! f(x)= -x+4

OpenStudy (anonymous):

ok so just plug 1 and -1 in for x: x = 1... -(1) + 4 = 3 x = -1... -(-1) +4 = 5 doesn't look like b. f(1) > f(-1) is right...

OpenStudy (anonymous):

since f(1) = 3 and f(-1) = 5 3 > 5 is a false statement.

OpenStudy (anonymous):

ahh! I get it! Thank you! Godbless :)

OpenStudy (anonymous):

No Problem! :D